OpenStudy (anonymous):

how do you solve this- indefinite integration of(sin(x-a)/sin(x+a))^1/2

OpenStudy (anonymous):

The whole thing is squarerooted, that makes sense haha. http://www.wolframalpha.com/input/?i=integrate+%28sin%28x-a%29%2Fsin%28x-a%29%29%5E%281%2F2%29

OpenStudy (perl):

batman, you have a x -a in the denominator

OpenStudy (anonymous):

ashgkasdhgasgkh

OpenStudy (anonymous):

Yeah so it's something crazy

OpenStudy (perl):

it doesnt looks like wolfram has a simple output http://www.wolframalpha.com/input/?i=integrate+%28sin%28x-a%29%2Fsin%28x%2Ba%29%29^%281%2F2%29

OpenStudy (perl):

woops i meant to say, double check your integral

OpenStudy (anonymous):

Nope, I'm not exactly sure how I would approach this

OpenStudy (perl):

its ok, i think we need more information from the original poster

ganeshie8 (ganeshie8):

wolfram is giving complex stuff because it is assuming x takes all real values

ganeshie8 (ganeshie8):

if it is a real valued function, we need to work it only when the ratio of sins inside the radical is nonnegative i think

OpenStudy (dan815):

x+a=/=npi, n E Z

OpenStudy (anonymous):

sin(x-a) = sin(x+a)cos2a-cos(x+a)sin2a sin(x+a) = cos2a-sin2a*cot(x+a) integrate idk

OpenStudy (dan815):

sin(A+B)=sin A cos B + cos A sin B

OpenStudy (dan815):

sin(A-B)==sin A cos -B + cos A sin -B =sinA cosB -CosA SinB

OpenStudy (dan815):

hmmm

OpenStudy (dan815):

now how about multipling by cojugate

ganeshie8 (ganeshie8):

if u want a feel for graph of function :

OpenStudy (dan815):

(sin A cos B + cos A sin B) (sin A cos B + cos A sin B) ---------------------------------------------- sin A cos B - cos A sin B(sin A cos B + cos A sin B)

OpenStudy (dan815):

is it possible to change this whole thing to complex form??

OpenStudy (dan815):

and take the imaginary part as solution

ganeshie8 (ganeshie8):

idk how we define area in complex

OpenStudy (dan815):

hey how about this thought

OpenStudy (dan815):

suppose i write sin(x-a) ------- sin(x+a) in complex form, now there shud be 4 seperate expansions and we want to

ganeshie8 (ganeshie8):

oh you want to integrate real and imaginary parts separately and add up ?

OpenStudy (dan815):

only consider the part where that complex expressio nis dealing with the one above

OpenStudy (dan815):

and we know it must exist in the real part, however the teal part will also include the cos(x-a)/cos(x+a)

ganeshie8 (ganeshie8):

i think we need to assume the given funciton is complex valued

OpenStudy (dan815):

the complex parst will be from cos(x-a) ------ -sin(x+a) and isin(x-a) -------- cos(x+a)

OpenStudy (dan815):

so we can rigt away eliminate the complex solution from our integral after we are done

OpenStudy (dan815):

since none of that contain the sin(x-a)/sin(x+a)

OpenStudy (dan815):

uhh nvm!

OpenStudy (dan815):

forgot there ssome real parts and im parts i dunnooo maybe ill work on it xD expand the whole thing and see if we can simplfy it somehow with complex

OpenStudy (dan815):

pretty sure its some trig sub stuff thoughhh

OpenStudy (anonymous):

i really tried solving this, and dan815, i tried expanding sin(x+a) and sin(x-a) and solving, but that only makes a bigger mess of things.

OpenStudy (anonymous):

the answer is:\[\cos a. arc \cos(\cos x \div \cos a)-\sin a.\ln(\sin x+\sqrt{\sin²x-\sin²a})+c\]